OpenStudy (anonymous):

Find the amount of the discount on a $190 bedspread that is now selling for 60% off. Then find the sale price of the bedspread.

OpenStudy (anonymous):

Amount of the discount = 0.4 * 190 = 76 Sale price = 190-76=114

OpenStudy (anonymous):

so its 114

OpenStudy (anonymous):

The sale price is 114 The discount is 76